OpenStudy (anonymous):

What is the discriminant of 4x^2 – 3x + 2 = 0? Answer A. 35 B. 41 C. –24 D. –23 What is the discriminant of 4x^2 – 3x + 2 = 0? Answer A. 35 B. 41 C. –24 D. –23 @Mathematics

OpenStudy (anonymous):

\[b^2-4ac\] with \[a=4,b=-3,c=2\] \[(-3)^2-4\times 4\times 2\] \[9-32\] \[-21\]

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

(-3)^2 - 4*4*2

OpenStudy (anonymous):

Lmfao

OpenStudy (anonymous):

maybe \[9-32=-23\]?

OpenStudy (hoblos):

-23

OpenStudy (anonymous):

yup -23

OpenStudy (anonymous):

Thanks, guys lol

OpenStudy (anonymous):

not my strong point (arithmetic)

OpenStudy (anonymous):

tut tut lol